OpenStudy (anonymous):

if d=1 and h= -2 what is the equation

OpenStudy (thejax):

ho ha where from!

OpenStudy (skullpatrol):

Those are two equations.

OpenStudy (anonymous):

d+h=-1 dh=-2 d/h=-1/2 h/d=-2